OpenStudy (tumblewolf):

Help with a few Calculus problems. WILL FAN AND MEDAL I need to find the derivitave of a couple problems and simple can't figure them out F (x)= 4x^4-(1/×^2) F (x)=(1/4-x^2) F (x)=(2x^3-3x^2+4x-5)/(x^2) U (x)=1/(x+2) ^2

jimthompson5910 (jim_thompson5910):

for the first one, it might help to think of 1/x^2 as x^(-2)

jimthompson5910 (jim_thompson5910):

You'll use the power rule if y = x^n then dy/dx = n*x^(n-1)

OpenStudy (tumblewolf):

Would the first one be 16X^3+2x?

jimthompson5910 (jim_thompson5910):

the derivative of x^(-2) is -2x^(-3) notice how I pulled down the exponent and then subtracted 1 from -2 to get -3

jimthompson5910 (jim_thompson5910):

-2x^(-3) = -2/(x^3)

OpenStudy (tumblewolf):

So 16x^3+2x^3

jimthompson5910 (jim_thompson5910):

more like \[\Large 16x^3 - \frac{2}{x^3}\]

OpenStudy (tumblewolf):

Ah ok now I see

jimthompson5910 (jim_thompson5910):

the second one isn't really clear, can you draw it out?

OpenStudy (tumblewolf):

No I'm on my tablet but it's 1 divided by 4-x^2 Is that any better? I'm sorry!

jimthompson5910 (jim_thompson5910):

ok so you'll need to use the quotient rule for this one

OpenStudy (tumblewolf):

That's (low)(d'high)-(high)(d'low) right?

jimthompson5910 (jim_thompson5910):

yeah I'm used to this formula if \[\Large h(x) = \frac{f(x)}{g(x)}\] then \[\Large h \ '(x) = \frac{f \ '(x)*g(x) - f(x)g \ '(x)}{[g(x)]^2}\]

OpenStudy (tumblewolf):

Yes that's it. I can't figure it all out though it's confusing me

jimthompson5910 (jim_thompson5910):

in this case f(x) = 1 g(x) = 4-x^2 what are `f ' (x)` and `g ' (x)` equal to?

OpenStudy (tumblewolf):

F (x)= 0 G (x)= 2x

OpenStudy (tumblewolf):

That's the part that throws me

jimthompson5910 (jim_thompson5910):

g ' (x) = -2x

OpenStudy (tumblewolf):

Oh right

jimthompson5910 (jim_thompson5910):

\[\Large h \ '(x) = \frac{f \ '(x)*g(x) - f(x)g \ '(x)}{[g(x)]^2}\] \[\Large h \ '(x) = \frac{0*(4-x^2) - 1*(-2x)}{(4-x^2)^2}\] now simplify

OpenStudy (tumblewolf):

2x/(4-x^2)^2

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

you'll use the quotient rule for `(2x^3-3x^2+4x-5)/(x^2)`

OpenStudy (tumblewolf):

So (X^2)(6x^2-6x+4)-(2x^3-3x^2+4x-5)(2x)

jimthompson5910 (jim_thompson5910):

that's all over [g(x)]^2 = [x^2]^2 = x^4

OpenStudy (tumblewolf):

Then I foil and combine what I can?

jimthompson5910 (jim_thompson5910):

yes simplify as much as possible

OpenStudy (tumblewolf):

(6x^4-6x^2+4x^2-4x^4-6x^3+8x^2-10x)

OpenStudy (tumblewolf):

10x^4-2x^2+6x^3-10x

jimthompson5910 (jim_thompson5910):

I'm not getting the same

OpenStudy (tumblewolf):

What am I doing wrong?

jimthompson5910 (jim_thompson5910):

(x^2)*(6x^2-6x+4)-(2x^3-3x^2+4x-5)*(2x) (x^2)*(6x^2)+(x^2)*(-6x)+(x^2)*(4)-2x*2x^3-2x*(-3x^2)-2x*4x-2x*(-5) 6x^4-6x^3+4x^2-4x^4+6x^3-8x^2+10x 6x^4-4x^4-6x^3+6x^3+4x^2-8x^2+10x 2x^4-4x^2+10x

jimthompson5910 (jim_thompson5910):

so you'll have 2x^4-4x^2+10x all over x^4

jimthompson5910 (jim_thompson5910):

then you can factor out x from the numerator and have it cancel with one factor of x from the denominator

OpenStudy (tumblewolf):

2x^3-4x+10

OpenStudy (tumblewolf):

Over x^3

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (tumblewolf):

It's always the factoring and foiling

OpenStudy (tumblewolf):

So we use the quotient rule again? (x+2)^2 (0)-(1)(2x)

OpenStudy (tumblewolf):

Is it just 2x/(x+4)^2

OpenStudy (mathmale):

Could you possibly post just one problem at a time? I do not readily recognize what your "Is it just 2x/(x+4)^2" refers to. A new problem? what are the instructions for it?

OpenStudy (tumblewolf):

Just find the derivitave. I was referring to the last problem on the last after it's been simplfied